OpenStudy (anonymous):

A track has lanes 1 meter wide. The turn radius of the inner lane is 25 meters. To make a fair race, the starting line in each lane must be staggered so each competitor runs teh same distance to the finish line. find the distanc between the starting line in one lane to the starting lane in the next lane. Is the same distance used between first and second lanes and between second and third lanes?

OpenStudy (anonymous):

this would depends on width of each lane. is there a specification for that?

OpenStudy (anonymous):

Assuming that there are N lanes and runners are running in the middle of each lane, width of each lane = (1/N) [m] turning of first lane = \[\pi \times (R + (1/N) \times (1/2))\] turning of second lane = \[\pi \times (R + ((1/N) \times (1/2)) + (1/N))\] you can get the difference in turning lane simply by subtracting turning of first lane from turning of second lane. If i worked the problem correctly, it should be \[\pi \times (1/N)\] meter difference between each lane